150.0 mL of 0.23 M  HF with 220.0 mL of 0.31 M  NaF

Calculate the pH of the solution that results from each of the following mixtures.

also, I'm so confused how you get the pKa..... can someone explain that part
Answer #1

pKa of HF=3.17

[HF]=(150*10^-3)*0.23=0.0345M

[NaF]=(220*10^-3)*0.31=0.0682M

pH=pKa+log[NaF]/[HF]=3.17+log(0.0682/0.0345)=3.46